# On the Challenges and Opportunities in Generative AI

### v3

@@ 2.4 Foundation Models for Heterogeneous Data Types
While there has been tremendous progress in large-scale foundation models for modalities like text and images, as the scope of application widens to encompass a broader range of data modalities, a variety of challenges emerge. These challenges are particularly pronounced in specialized fields such as healthcare and chemistry (Raghupathi & Raghupathi, 2014; Korshunova et al., 2022).
@@
In healthcare, generation based on diverse data types—including imaging, health records, and genomics—poses challenges in interoperability, data privacy, and security (Moor et al., 2023). Time series generation, in particular, requires addressing irregularly sampled data, missing values, seasonality, and long-term dependencies (Steinberg et al., 2021). In chemistry, physics, and chemical engineering, generative models have huge potential, not just for molecule, drug, and material design, but also in data augmentation, property prediction, and reaction prediction (Winter et al., 2019; Ahmad et al., 2022; Castro Nascimento & Pimentel, 2023; Hu et al., 2020). Data in these fields are often sparse, heterogeneous and correlated. On the other hand, they provide a vast body of physical and chemical domain knowledge, ranging from (strict) laws of nature and boundary conditions to (soft) empirical correlations and human experience. Therefore, developing hybrid (ML + domain knowledge) foundation models is a particular challenge (Venkatasubramanian, 2019; Jirasek et al., 2022; Jirasek & Hasse, 2023; Howard et al., 2022).
The overarching goal is to build general models, which can be applied to different applications spanning various data type, by integrating information from diverse sources and understanding complex relationships across different types of data (Li et al., 2023a; Reed et al., 2022; Driess et al., 2023). However, to effectively interact with the physical world, embodied agents must integrate perception, reasoning, and planning across vision and language, along with executing physical actions and environmental interactions. While datasets for natural language or images are relatively accessible (Hausknecht et al., 2020; Li et al., 2023b; 2024b), a comparable dataset for control tasks is lacking, suggesting generative simulation as a potential solution (Xian et al., 2023; Fan et al., 2022).
@@ 3 Optimizing Efficiency and Resource Utilization
Efforts to scale deep generative models (DGMs) for tasks like language modeling and text-to-image synthesis often involve training large models with billions of parameters, which demands significant computational resources. This leads to practical issues such as high energy costs (Wu et al., 2022) and expensive inference, limiting access for many users. There is a clear need to reduce the memory and computational requirements of large-scale DGMs to enhance accessibility and sustainability (Bender et al., 2021).
@@ 3.1 Efficient Training and Inference
Optimizing the network architecture, which forms the backbone of modern machine learning, is crucial for efficient training and inference in DGMs. While we have seen recent improvements in model quality (OpenAI, 2023; Touvron et al., 2023; Peebles & Xie, 2023), a principled investigation of the role of different underlying neural network modules is still lacking. For instance, several popular LLMs like PaLM (Chowdhery et al., 2022) and Llama (Touvron et al., 2023) still largely reuse the original transformer architecture from Vaswani et al. (2017) with some additional modifications (Shazeer, 2020; Su et al., 2024; Zhang & Sennrich, 2019). A modification of particular importance has been that of the self-attention (Bahdanau et al., 2015) mechanism; in the original architecture, this operation incurred a computational cost that scaled quadratically in the context length. This made inference computationally expensive, especially for long-context modeling. Several recent works have proposed attention variants that provide faster inference times (Tay et al., 2022). For example, Flash Attention (Dao et al., 2022) employs hardware optimizations and efficient memory management techniques to reduce the effective computational overhead of attention from quadratic to linear in the context length. Grouped Query Attention (Ainslie et al., 2023) proposes a structural change to the standard attention mechanism, where queries [^4] A query can loosely be thought of as a transformed vector representation of a token from the input sequence. are divided into distinct groups that are then processed independently and simultaneously. Methods that make these popular network architectures more computationally efficient—as well as alternative autoregressive sequence-modeling frameworks with favorable properties like scalability and linear complexity in the context length (Gu & Dao, 2023; Gu et al., 2021)—remain an interesting direction for future work.
@@
Despite some progress in developing PTQ and QAT methods for LLMs (Dettmers et al., 2022; Liu et al., 2023b; Xiao et al., 2023; Yao et al., 2022) and large-scale text-to-image diffusion models (Li et al., 2023c), the existing methods are far from perfect. For instance, OPTQ (Frantar et al., 2023), a PTQ-based approach, can perform inference for a quantized LLM (in this case OPT (Zhang et al., 2022)) with 175B parameters on a single A100 GPU with 80GB of memory without degradation in accuracy. Though impressive, even this quantized model would likely have limited utility on a consumer-grade GPU device, let alone on standard edge devices. Similarly, QAT-based approaches can often achieve lower bitrates but trade off additional training for this efficiency. This can be a major computational bottleneck for large generative models. While some recent work suggests preliminary success in this direction (Lin et al., 2024a), we believe that investigating the impact of model quantization at low bitrates in large-scale generative models is a crucial direction for the practical deployment of these models.
@@
To address the inefficiencies in diffusion models, researchers have explored multiple complementary approaches to speed up inference. Some notable approaches include: developing training-free samplers (Song et al., 2021; Liu et al., 2022a; Lu et al., 2022; Zhang & Chen, 2023; Karras et al., 2022; Pandey et al., 2024), designing better diffusion processes (Singhal et al., 2023; Dockhorn et al., 2022; Pandey & Mandt, 2023; Karras et al., 2022), and combining other model families with diffusion models (Pandey et al., 2022; Zheng et al., 2023b; Xiao et al., 2022; Yang & Mandt, 2023). Additionally, training a diffusion model in the latent space of a lossy transform (Vahdat et al., 2021; Rombach et al., 2022) not only improves memory requirements and sampling efficiency but also provides access to a more interpretable low-dimensional latent representation. A lossy transform (such as VQ-GAN (Esser et al., 2021)) can drastically reduce data dimensionality while retaining the perceptually relevant details of high-resolution images. Designing more efficient lossy compression operations in the context of diffusion models has received less attention in the community and is an important direction for further work (Yang et al., 2023b; Havasi et al., 2019; Yang et al., 2020). Despite these advances, sampling from diffusion models remains computationally challenging, typically requiring 25-50 NFEs to generate high-quality samples. While approaches based on progressive distillation (Salimans & Ho, 2022; Meng et al., 2023) can further speed up inference, they trade off additional training for faster sampling. Therefore, there is a need for DGMs that inherit all the advantages of diffusion models while supporting one-step sample generation by design (e.g., see consistency models (Song et al., 2023; Song & Dhariwal, 2024) for recent work in this direction).
@@
In the case of LLMs, in addition to an expensive self-attention operation in transformer-based autoregressive models, sequential token generation in a left-to-right fashion in these models makes inference more expensive. Indeed, it is one reason that (sub)word tokenization—the pre-processing of text into pre-defined units—is still an essential part of these pipelines. Notably, tokenization itself introduces a strong inductive bias into language modeling: the model is constrained to work with the predefined units set by the tokenization scheme. The representation of the data that the model learns is inherently shaped by these units, limiting the model’s flexibility. Token-free approaches have been proposed to allow for the joint optimization of text segmentation alongside other parameters, but in practice, they are often computationally infeasible with attention-based architectures because handling raw character sequences at scale magnifies the already expensive attention mechanism (Xue et al., 2022). Consequently, tokenization remains a key design choice that shapes both the performance and efficiency of modern generative models and whose further optimization is constrained by needs for efficiency (Rust et al., 2021; Toraman et al., 2023; Ali et al., 2024). Methods such as speculative decoding (Leviathan et al., 2022) are one approach that can help reduce the bottleneck caused by left-to-right generation. Non-autoregressive models also offer an interesting alternative for sequence modeling. For example, diffusion models amortize the computational cost of generating sequences across all tokens simultaneously (Dieleman et al., 2022; Wu et al., 2023; Li et al., 2022a). However, these models inherently lack the inductive bias for contextual generation, which has been shown to work well empirically for sequential modeling tasks. This affects their performance in downstream tasks that might require long-context modeling, such as video synthesis (Yang et al., 2023a; Ho et al., 2022a). While diffusion models can be incorporated within the autoregressive framework for such tasks, the resulting models can be very expensive during inference (due to the cost of synthesizing a single token using diffusion across multiple tokens). Therefore, we identify a potential tradeoff between long context modeling and efficient inference, with the diffusion and autoregressive modeling paradigms falling on the opposite ends of this tradeoff. Hence, designing generative modeling paradigms that can optimally balance this tradeoff remains challenging.
@@ 3.2 Evaluation Metrics
Many generative models are probabilistic, making likelihood-based metrics a seemingly natural choice for evaluating their performance. These metrics have been widely utilized in the literature due to their alignment with the probabilistic frameworks of such models. However, empirical evidence suggests that likelihood-based metrics often do not provide an accurate assessment of generation quality (Theis et al., 2016). In particular, they often fail to correlate with human judgments of sample quality (Kolchinski et al., 2019; Jiralerspong et al., 2023; Pimentel et al., 2023). Moreover, many popular generative models do not even allow for tractable likelihood computation. Other automatic evaluation metrics are thus necessary for evaluating the quality of generated samples.
@@
Several notable evaluation metrics for generative models follow the general paradigm of comparing the distribution of generated samples to that of train/test samples (Sajjadi et al., 2018; Pillutla et al., 2024). For instance, the Fréchet inception distance (Heusel et al., 2017), which is widely used for evaluating image synthesis models, takes this approach (Salimans et al., 2016; Bińkowski et al., 2018). However, these metrics are far from perfect. First, robust computations of these metrics require a large set of samples (around 50k for image generation models). This can be computationally demanding for generative models with a sequential inference process, like diffusion and autoregressive models. Even given large sample sizes, these metrics have still demonstrated issues with robustness. For example, FID can be sensitive to minor perturbations in the input data (Parmar et al., 2022) (see Chong & Forsyth (2020) for additional discussion on sources of bias associated with FID and Borji (2022) for more related evaluation metrics). Second, these methods typically rely on an external pretrained model, e.g., the GPT-2 family of language models (Radford et al., 2019) or a classifier network trained on ImageNet (Deng et al., 2009). This property makes the metric effective for evaluating sample quality within the domain of the pretrained model’s data but seemingly causes it to overlook significant features or overemphasize arbitrary ones in other domains (Kynkäänniemi et al., 2023; Pimentel et al., 2023).
@@
Recent works have attempted to improve upon the above-mentioned shortcomings. For example, Jayasumana et al. (2024) propose the use of embeddings from the CLIP model (Radford et al., 2021), which aligns images and text in a shared embedding space, in order to make a more robust evaluation metric for image synthesis models. Several evaluation metrics for text generation systems prompt LLMs to score or rank samples (Li et al., 2024c). These metrics correlate remarkably well with human judgments (Kocmi & Federmann, 2023) and offer more fine-grained assessments of text generation systems—providing feedback at the individual sample level and taking into account user-specified criteria (Jiang et al., 2024a). These methods make progress towards broader applicability across domains and greater alignment with human evaluations but also demonstrate an increased reliance on generative models for the evaluation of other generative models. This circular dependency introduces the risk of amplifying existing model biases (Fang et al., 2024) and narrowing the diversity of model-generated content (Doshi & Hauser, 2024; Gambetta et al., 2024), as the underlying paradigm of such evaluation metrics should lead to the favoring of outputs that align with the characteristics and biases of the models used for assessment. Some works have proposed explicitly rewarding sample diversity in evaluation metrics (Zhu et al., 2018; Alihosseini et al., 2019), which would alleviate the latter problem. However, there is often a quality-diversity trade-off (Caccia et al., 2019; Zhang et al., 2021; Naeem et al., 2020), where a model that generates high-quality samples might have low diversity across its samples and vice versa. Further, the quantification of diversity is in itself a difficult task (Tevet & Berant, 2021).
@@
Subjective aspects in generation. A major challenge underlying the evaluation of generation quality is the subjective nature of sample attributes, such as realism, fluency, and style. While human inspection is typically the gold standard for evaluating generative models (Denton et al., 2015; Zhou et al., 2019; Saharia et al., 2022), in many cases, human judges disagree over several attributes, such as which samples have better quality (Clark et al., 2021) or what is considered realistic in the target domain (e.g., medical images or industrial optical inspection). This challenge is even present for conditional synthesis tasks when the set of suitable outputs is limited by constraints from the input. For example, in text-to-image generation (Ramesh et al., 2021; Rombach et al., 2022; Saharia et al., 2022), human evaluators may have different opinions on how closely a generated image aligns with the provided description. .
@@
Evaluating Model Uncertainty and Calibration. Model uncertainty and calibration have become quantities of interest because of their implications for the reliability, interpretability, and safety of generative AI systems. Here, we use the term model uncertainty to refer to the degree of confidence a model has in its outputs; model calibration refers to the degree to which a model’s estimated probability of an event is consistent with that event’s true probability of occurring. [^6] We note that other definitions for the terms have been used. As concrete examples of the importance of these metrics, high model uncertainty in language generation tasks has been linked with the occurrence of hallucinations—instances where the model produces outputs that are implausible or factually incorrect (van der Poel et al., 2022; Zhang et al., 2023b); autonomous driving systems increasingly use generative models to predict the trajectories of other vehicles, cyclists, and pedestrians (Yuan et al., 2020), and miscalibration of the probabilities of such trajectories can lead to severe accidents.
@@
Model Selection. Model selection, i.e., identifying which model configuration or set of hyperparameters will perform best on a given task, is essential in training large-scale generative models. Evaluation metrics play a critical role here. Using knowledge of scaling laws (Kaplan et al., 2020; Henighan et al., 2020), evaluation metrics can be used to predict early on in a training run whether a model is likely to be successful (OpenAI, 2023), potentially reducing the need to train numerous large neural networks in the search for a single good model. We also believe that more effort should be invested in analyzing models’ performance-complexity tradeoff, an important yet under-investigated measure for real-world applications at scale. This tradeoff refers to the balance between model performance and computational complexity. We argue that model selection and evaluation should perhaps shift towards identifying the model families that lie in the associated Pareto set that optimizes this tradeoff (Devroye, 2010; Braverman, 2005; Chen et al., 2022; Braverman, 2023), as optimizing for these characteristics in isolation does not account for real-world constraints. The naïve approach—training well-performing models in each of the model classes under consideration and computing their respective computational complexities—is time- and resource-intensive. We posit that alternative assessments of complexity from information and learning theory (Xu et al., 2020, e.g.,) could provide more efficient substitutes for these types of evaluations.
@@
Evaluation metrics can help us understand and ultimately mitigate model shortcomings. While this approach has been embraced for improving model quality (Ranzato et al., 2016; Black et al., 2024, inter alia), it also has significant potential to enhance model fairness, safety, and reliability. For instance, metrics designed to quantify various forms of bias can aid in identifying and addressing model unfairness. While such metrics exist for classification or regression models (e.g., demographic parity or equalized odds), their extension to generative models is non-trivial. Research is thus needed to develop and refine metrics that can effectively quantify biases in the complex outputs of generative models. [^7] Many aspects of fairness cannot be captured by quantitative metrics. Further, definitions of fairness can differ amongst different people and groups, and these definitions may evolve over time. However, they can still provide insights into whether models achieve a certain level of fairness in specific aspects. This includes creating frameworks that account for the nuanced and context-dependent nature of generated content, ensuring these models are not only high-quality but also fair and aligned with ethical standards (Ray, 2023). Unfortunately, to be effective, these metrics must also be adaptable to closed-source generative models since parameters and logits of most commercial models are not publicly available (Zhao et al., 2023a; Sun et al., 2023; Laszkiewicz et al., 2024).

### v4

@@ 2.4 Foundation Models for Domain-specific and Heterogeneous Modalities
While there has been tremendous progress in large-scale foundation models for modalities like text and images, as the scope of their application widens to encompass a broader range of data modalities, a variety of challenges surrounding cross-modal alignment, data interoperability, privacy, and evaluation emerge. These challenges are particularly pronounced in specialized fields such as healthcare and chemistry (Raghupathi & Raghupathi, 2014; Korshunova et al., 2022; Busch et al., 2025).
@@
In healthcare, generation based on diverse data types—including imaging, health records, and genomics—poses challenges in interoperability, data privacy, and security (Moor et al., 2023a). Time series generation, in particular, requires addressing irregularly sampled data, missing values, seasonality, and long-term dependencies (Steinberg et al., 2021). In chemistry, physics, and chemical engineering, generative models have huge potential, not just for molecule, drug, and material design, but also in data augmentation, property prediction, and reaction prediction (Winter et al., 2019; Ahmad et al., 2022; Castro Nascimento & Pimentel, 2023; Hu et al., 2020). Data in these fields are often sparse, heterogeneous and correlated. On the other hand, they provide a vast body of physical and chemical domain knowledge, ranging from (strict) laws of nature and boundary conditions to (soft) empirical correlations and human experience. Therefore, developing hybrid (ML + domain knowledge) foundation models is a particular challenge. While there has been some recent progress toward this goal, e.g., in the realms of physics (Jirasek et al., 2022; Jirasek & Hasse, 2023; Howard et al., 2022) and medicine (Raghu et al., 2022; Moor et al., 2023b; Xia et al., 2024), there is still much work to be done (Venkatasubramanian, 2019).
We argue that an overarching goal of the generative modeling field is to build general models that can seamlessly integrate information from diverse sources and understand complex relationships across different types of data (Li et al., 2023a; Reed et al., 2022; Driess et al., 2023). This multi-modal integration challenge becomes particularly evident in embodied AI applications, where generative models must serve multiple functions simultaneously. We can thus view embodied agents as a natural testing ground for truly general generative AI because of their broad requirements: (1) generative world modeling to predict future states and simulate outcomes of potential actions, (2) multi-modal generation to produce coherent plans that span language instructions, visual predictions, and motor commands, and (3) conditional generation that respects physical constraints while taking into account dynamics of the environment. Unlike purely digital applications, where errors may be aesthetic or semantic, embodied systems expose fundamental limitations in our generative models through physical failure, e.g., a robot failing to grasp objects or navigate spaces. However, developing these systems faces a significant hurdle: data scarcity. While datasets for natural language or images are relatively accessible (Hausknecht et al., 2020; Li et al., 2023b; 2024f), multi-task, multi-environment datasets of control trajectories (states, actions, and outcomes) remain comparatively rare, hindering progress on the development of generalist embodied agents. Generative simulation is one route we identify to achieve this potential use case for cross-domain generative models (Xian et al., 2023; Fan et al., 2022).
@@ 3 Optimizing Efficiency and Resource Utilization
Sub-challenge Typical failure mode Promising mitigation avenues Reference materials Research-question candidates
Efficient attention mechanisms (§3.1) Long contexts needed in certain settings and context length limited by computational demands Hardware-aware attention computations (e.g., FlashAttention-2); Sub-quadratic attention alternatives Datasets:
LongBench
Surveys:
Tay et al. (2022a)
Q1: Can (non-autoregressive) selective-state SSMs (e.g. Mamba) achieve the same performance as transformers in text generation tasks?
Q2: Can better retrieval methods in RAG systems mitigate the need for longer context windows in LLMs?
Low-bit quantization without quality loss (§3.1) Sharp accuracy drop when reducing FP precision to less than 4-bits Activation-aware weight quantization; Quantization-aware training and fine-tuning Surveys:
Zeng et al. (2025)
Q1: Do activation-aware quantization methods preserve model calibration after preference-based fine-tuning (e.g., RLHF)?
Q2: What theoretical limits bound post-training quantization of diffusion models?
Fast sampling for diffusion models (§3.1) Hundreds of network evaluations needed per sample Progressive Distillation; Consistency Models; Model quantization Benchmarks:
FID/IS on CIFAR-10/ImageNet
Surveys:
Shen et al. (2025)
Q1: Can we design models that achieve one-step generation while maintaining diffusion models’ training stability and sample quality?
Reliable quality metrics (§3.2) Automatic evaluation metrics do not correlate with human perception of quality Generative models for quality assessment (e.g., LLM-as-Judge, Auto-J); Sample-based metrics (e.g., Feature-Likelihood Score) Benchmarks:
JudgeBench
Surveys:
(Betzalel et al., 2024)
Q1: How can learned reward models be made reproducible enough to serve as public benchmarks?
Q2: Is a unified multi-modal mauve variant feasible?
Q3: How reliable are LLM-as-Judge scores in OOD settings?
Compute-efficient model selection (§3.2) Brute force grid-search approaches to model selection are prohibitive for today’s large models Scaling-law extrapolation; Zero-cost proxies Benchmarks:
NAS-Bench-101
; NATS-Bench
Surveys:
Li et al. (2024a)
White et al. (2022)
Q1: Can information-theoretic complexity measures be used during model architecture search to reliably rule out entire classes of models?
Q2: Can zero-cost proxies and scaling law extrapolations be effectively combined to provide a stronger indication of optimal models than their individual signals?
Table 2: Research challenges summary table - Optimizing Efficiency and Resource Utilization
Efforts to scale DGMs for tasks like language modeling and text-to-image synthesis often involve training large models with billions of parameters, which demands significant computational resources. This leads to practical issues such as high energy costs (Wu et al., 2022) and expensive inference, limiting access for many users. This further raises environmental concerns due to the energy consumption required to fuel modern tensor processing hardware (Strubell et al., 2020). Training PaLM leads to $271$ tons of CO2e effective emissions (Chowdhery et al., 2022) and training GPT-3 emits roughly twice as much under comparable accounting assumptions (Patterson et al., 2022). Therefore, there is a clear need to reduce the memory and computational requirements of large-scale DGMs to enhance accessibility and sustainability (Bender et al., 2021).
@@ 3.1 Efficient Training and Inference
Optimizing the network architecture, which forms the backbone of modern machine learning, is crucial for efficient training and inference in DGMs. While we have seen recent improvements in model quality (OpenAI, 2023; Touvron et al., 2023; Peebles & Xie, 2023), there is still a dearth of systematic comparative studies of architectural components’ contributions to generative model performance. For instance, several popular LLMs like PaLM (Chowdhery et al., 2022) and Llama (Touvron et al., 2023) still largely reuse the original transformer architecture from Vaswani et al. (2017) with some additional modifications (Shazeer, 2020; Su et al., 2024; Zhang & Sennrich, 2019). A modification of particular importance has been that of the self-attention (Bahdanau et al., 2015) mechanism; in the original architecture, this operation incurred a computational cost that scaled quadratically in the context length. This made inference computationally expensive, especially for long-context modeling. Several recent works have proposed attention variants that provide faster inference times (Tay et al., 2022b). For example, Flash Attention (Dao et al., 2022) employs hardware optimizations and efficient memory management techniques to reduce the effective computational overhead of attention from quadratic to linear in the context length; Flash Attention 2 (Dao, 2024) takes these optimizations a step further, bringing attention computations close to the achievable bounds on fast matrix multiplication. Grouped Query Attention (Ainslie et al., 2023) proposes a structural change to the standard attention mechanism, where queries [^5] In the attention mechanism, a query is a transformed vector representation, typically derived from input tokens. are divided into distinct groups that are then processed independently and simultaneously. We see several promising research directions for reducing the computational needs of large-scale generative models, including specialized methods that make popular network architectures more computationally efficient (e.g., the attention variants discussed above), early-exit designs that allow models to make predictions without running the forward pass through the full network (Chen et al., 2024b) and alternative autoregressive sequence-modeling frameworks with favorable properties like scalability and linear complexity in the context length (Gu & Dao, 2023; Gu et al., 2021).
@@
Despite some progress in developing PTQ and QAT methods for LLMs (Dettmers et al., 2022; Liu et al., 2023b; Xiao et al., 2023; Yao et al., 2022; Dettmers et al., 2023) and large-scale text-to-image diffusion models (Li et al., 2023d), the existing methods are far from perfect. For instance, OPTQ (Frantar et al., 2023), a PTQ-based approach, can perform inference for a quantized LLM (in this case OPT (Zhang et al., 2022)) with 175B parameters on a single A100 GPU with 80GB of memory without degradation in accuracy. Though impressive, even this quantized model would likely have limited utility on a consumer-grade GPU device, let alone on standard edge devices. Similarly, QAT-based approaches can often achieve lower bitrates but trade off additional training for this efficiency. This can be a major computational bottleneck for large generative models. While some recent work suggests preliminary success in this direction (Lin et al., 2024a), we believe that investigating the impact of model quantization at low bitrates in large-scale generative models is a crucial direction for the practical deployment of these models.
@@
To address the inefficiencies in diffusion models, researchers have explored multiple complementary approaches to speed up inference. Some notable approaches include: developing training-free samplers (Song et al., 2021; Liu et al., 2022a; Lu et al., 2022; Zhang & Chen, 2023; Karras et al., 2022; Pandey et al., 2024), designing better diffusion processes (Singhal et al., 2023; Dockhorn et al., 2022; Pandey & Mandt, 2023; Karras et al., 2022), and combining other model families with diffusion models (Pandey et al., 2022; Zheng et al., 2023b; Xiao et al., 2022; Yang & Mandt, 2023). Additionally, training a diffusion model in the latent space of a lossy transform (Vahdat et al., 2021; Rombach et al., 2022) not only improves memory requirements and sampling efficiency but also provides access to a more interpretable low-dimensional latent representation. A lossy transform (such as VQ-GAN (Esser et al., 2021)) can drastically reduce data dimensionality while retaining the perceptually relevant details of high-resolution images. Designing more efficient lossy compression operations in the context of diffusion models has received less attention in the community and is an important direction for further work (Yang et al., 2023d; Havasi et al., 2019; Yang et al., 2020). Despite these advances, sampling from diffusion models remains computationally challenging, typically requiring 25-50 NFEs to generate high-quality samples. While approaches based on progressive distillation (Salimans & Ho, 2022; Meng et al., 2023) can further speed up inference, they trade off additional training for faster sampling. Therefore, there is a need for DGMs that inherit all the advantages of diffusion models while supporting one-step sample generation by design (e.g., see consistency models (Song et al., 2023; Song & Dhariwal, 2024) for recent work in this direction).
@@
In the case of LLMs, in addition to an expensive self-attention operation in transformer-based autoregressive models, sequential token generation in a left-to-right fashion in these models makes inference more expensive. Indeed, it is one reason that (sub)word tokenization—the pre-processing of text into pre-defined units—is still an essential part of these pipelines. Notably, tokenization itself introduces a strong inductive bias into language modeling: the model is constrained to work with the predefined units set by the tokenization scheme. The representation of the data that the model learns is inherently shaped by these units, limiting the model’s flexibility. Token-free approaches have been proposed to allow for the joint optimization of text segmentation alongside other parameters, but in practice, they are often computationally infeasible with attention-based architectures because handling raw character sequences at scale magnifies the already expensive attention mechanism (Xue et al., 2022). Dynamic tokenization schemes (Ahia et al., 2024, e.g.,) present an interesting research direction, as they allow for predicting token boundaries at inference time, but they likewise can introduce significant computational overhead. Tokenization remains a key design choice that shapes both the performance and efficiency of modern generative models and whose further optimization is constrained by needs for efficiency (Rust et al., 2021; Toraman et al., 2023; Ali et al., 2024). Methods such as speculative decoding (Leviathan et al., 2022) are one approach that can help reduce the bottleneck caused by left-to-right generation. Non-autoregressive models also offer an interesting alternative for sequence modeling. For example, diffusion models amortize the computational cost of generating sequences across all tokens simultaneously (Dieleman et al., 2022; Wu et al., 2023; Li et al., 2022a). However, these models inherently lack the inductive bias for contextual generation, which has been shown to work well empirically for sequential modeling tasks. This affects their performance in downstream tasks that might require long-context modeling, such as video synthesis (Yang et al., 2023c; Ho et al., 2022a). While diffusion models can be incorporated within the autoregressive framework for such tasks, the resulting models can be very expensive during inference (due to the cost of synthesizing a single token using diffusion across multiple tokens). Therefore, we identify a potential tradeoff between long context modeling and efficient inference, with the diffusion and autoregressive modeling paradigms falling on the opposite ends of this tradeoff. Hence, designing generative modeling paradigms that can optimally balance this tradeoff remains challenging.
@@ 3.2 Evaluation Metrics
Many generative models are probabilistic, making likelihood-based metrics a seemingly natural choice for evaluating their performance. These metrics have been widely utilized in the literature due to their alignment with the probabilistic frameworks of such models. However, empirical evidence suggests that likelihood-based metrics often do not provide an accurate assessment of generation quality (Theis et al., 2016). In particular, they often fail to correlate with human judgments of sample quality (Kolchinski et al., 2019; Pimentel et al., 2023). Moreover, many popular generative models do not even allow for tractable likelihood computation. Other automatic evaluation metrics are thus necessary for evaluating the quality of generated samples.
@@
Several notable evaluation metrics for generative models follow the general paradigm of comparing the distribution of generated samples to that of train/test samples (Sajjadi et al., 2018; Pillutla et al., 2024; Jiralerspong et al., 2023). For instance, the Fréchet inception distance (Heusel et al., 2017), which is widely used for evaluating image synthesis models, takes this approach (Salimans et al., 2016; Bińkowski et al., 2018). However, these metrics are far from perfect. First, robust computations of these metrics require a large set of samples (around 50k for image generation models). This can be computationally demanding for generative models with a sequential inference process, like diffusion and autoregressive models. Even given large sample sizes, these metrics have still demonstrated issues with robustness. For example, FID can be sensitive to minor perturbations in the input data (Parmar et al., 2022) (see Chong & Forsyth (2020) for additional discussion on sources of bias associated with FID and Borji (2022) for more related evaluation metrics). Second, these methods typically rely on an external pretrained model, e.g., the GPT-2 family of language models (Radford et al., 2019) or a classifier network trained on ImageNet (Deng et al., 2009). This property makes the metric effective for evaluating sample quality within the domain of the pretrained model’s data but seemingly causes it to overlook significant features or overemphasize arbitrary ones in other domains (Kynkäänniemi et al., 2023; Pimentel et al., 2023).
@@
Recent works have attempted to improve upon the above-mentioned shortcomings. For example, Jayasumana et al. (2024) propose the use of embeddings from the CLIP model (Radford et al., 2021), which aligns images and text in a shared embedding space, in order to make a more robust evaluation metric for image synthesis models. Several evaluation metrics for text generation systems use LLMs to score or rank samples, either via prompting (Li et al., 2024g) or explicit fine-tuning on the task of evaluation (Li et al., 2024e). These metrics correlate remarkably well with human judgments (Kocmi & Federmann, 2023) and offer more fine-grained assessments of text generation systems—providing feedback at the individual sample level and taking into account user-specified criteria (Jiang et al., 2024a). These methods make progress towards broader applicability across domains and greater alignment with human evaluations but also demonstrate an increased reliance on generative models for the evaluation of other generative models. This circular dependency introduces the risk of amplifying existing model biases (Fang et al., 2024) and narrowing the diversity of model-generated content (Doshi & Hauser, 2024; Gambetta et al., 2024), as the underlying paradigm of such evaluation metrics should lead to the favoring of outputs that align with the characteristics and biases of the models used for assessment. Some works have proposed (either explicitly or implicitly) rewarding sample diversity in evaluation metrics (Zhu et al., 2018; Alihosseini et al., 2019; Jiralerspong et al., 2023), which would alleviate the latter problem. However, there is often a quality-diversity trade-off (Caccia et al., 2019; Zhang et al., 2021; Naeem et al., 2020), where a model that generates high-quality samples might have low diversity across its samples and vice versa. Further, the quantification of diversity is in itself a difficult task (Tevet & Berant, 2021).
@@
Subjective aspects in generation. A major challenge underlying the evaluation of generation quality is the subjective nature of sample attributes, such as realism, fluency, and style. While human inspection is typically the gold standard for evaluating generative models (Denton et al., 2015; Zhou et al., 2019; Saharia et al., 2022), in many cases, human judges disagree over several attributes, such as which samples have better quality (Clark et al., 2021) or what is considered realistic in the target domain (e.g., medical images or industrial optical inspection). This challenge is even present for conditional synthesis tasks when the set of suitable outputs is limited by constraints from the input. For example, in text-to-image generation (Ramesh et al., 2021; Rombach et al., 2022; Saharia et al., 2022), human evaluators may have different opinions on how closely a generated image aligns with the provided description.
@@
Evaluating Model Uncertainty and Calibration. Model uncertainty and calibration have become quantities of interest because of their implications for the reliability, interpretability, and safety of generative AI systems. Here, we use the term model uncertainty to refer to the degree of confidence a model has in its outputs; model calibration refers to the degree to which a model’s estimated probability of an event is consistent with that event’s true probability of occurring. [^7] We note that other definitions for the terms have been used; we employ this one as it is the most relevant for our exposition. As concrete examples of the importance of these metrics, high model uncertainty in language generation tasks has been linked with the occurrence of hallucinations—instances where the model produces outputs that are implausible or factually incorrect (van der Poel et al., 2022; Zhang et al., 2023c); autonomous driving systems increasingly use generative models to predict the trajectories of other vehicles, cyclists, and pedestrians (Yuan et al., 2020), and miscalibration of the probabilities of such trajectories can lead to severe accidents.
@@
Model Selection. Model selection, i.e., identifying which model configuration or set of hyperparameters will perform best on a given task, is essential in training large-scale generative models. Evaluation metrics play a critical role here. Using knowledge of scaling laws (Kaplan et al., 2020; Henighan et al., 2020), evaluation metrics can be used to predict early on in a training run whether a model is likely to be successful (OpenAI, 2023). Recent work has shown that these predictions can be done quite precisely (Ruan et al., 2024), potentially reducing the need to train numerous large neural networks in the search for a single good model. Zero-shot proxies (Abdelfattah et al., 2021)—metrics computed on an untrained or minimally initialized network that approximate its final task performance before any training is done—are another promising research direction for compute-efficient model selection. We also believe that more effort should be invested in analyzing models’ performance-complexity tradeoff, an important yet under-investigated measure for real-world applications at scale. This tradeoff refers to the balance between model performance and computational complexity. We argue that model selection and evaluation should perhaps shift towards identifying the model families that lie in the associated Pareto set that optimizes this tradeoff (Devroye, 2010; Braverman, 2005; Chen et al., 2022; Braverman, 2023), as optimizing for these characteristics in isolation does not account for real-world constraints. The naïve approach—training well-performing models in each of the model classes under consideration and computing their respective computational complexities—is time- and resource-intensive. We posit that alternative assessments of complexity from information and learning theory (e.g., Xu et al., 2020) could provide the basis for more efficient metrics in these types of evaluations.
@@
Evaluation metrics can help us understand and ultimately mitigate model shortcomings. While this approach has been embraced for improving model quality, it also has significant potential to enhance model fairness, safety, and reliability. For instance, metrics designed to quantify various forms of bias can aid in identifying and addressing model unfairness. While such metrics exist for classification or regression models (e.g., demographic parity or equalized odds), their extension to generative models is non-trivial. Research is thus needed to develop and refine metrics that can effectively quantify biases in the complex outputs of generative models. [^8] Many aspects of fairness cannot be captured by quantitative metrics. Further, definitions of fairness can differ amongst different people and groups, and these definitions may evolve over time. However, they can still provide insights into whether models achieve a certain level of fairness in specific aspects. This includes creating frameworks that account for the nuanced and context-dependent nature of generated content, ensuring these models are not only high-quality but also fair and aligned with ethical standards (Ray, 2023). Unfortunately, to be effective, these metrics must also be adaptable to closed-source generative models since parameters and logits of most commercial models are not publicly available (Zhao et al., 2023a; Sun et al., 2023a; Laszkiewicz et al., 2024).